### v1

@@ 7 Limitations
While MiST demonstrates that targeted mid-stage pre-training can unlock chemical reasoning in a 3B-parameter model, several caveats remain. First, the SCS criterion works especially well in domains and tasks where the outcome validity is verifiable and easily corruptible, as is the case with chemistry and SMILES notation. We find such complications in the case of CMG, where CIF files are less-trivially corruptible, leading to worsened predictive power. Using a 100% valid notation, as is the case of SELFIES [83] for molecule generation tasks, can also compromise the value of SCS-like measures. Similar metrics can however be devised for biological sequential data, mathematical notation, etc. Second, the RLSF rewards we use focus on syntactic agreement with ground truth (e.g. exact SMILES or high Tanimoto similarity) and thus do not discourage chemically implausible or unsafe outputs, leaving open the possibility of reward hacking. Third, our evaluation suite—reaction prediction, IUPAC to SMILES translation, and conditional material generation, cover a narrow slice of chemistry; tasks that hinge on stereochemistry, kinetics, spectroscopy, or three-dimensional conformations remain unexplored. Finally, our pre-training corpus is dominated by small-molecule, chemical literature and patents, potentially biasing the model against macromolecular or bio-chemical domains. Addressing these limitations will be critical before SCS can be routinely used as a diagnostic metrics in other domains, and for MiST-style models to be relied upon as general scientific reasoning models.
@@ Acknowledgments
This work was supported as part of the Swiss AI Initiative by a grant from the Swiss National Supercomputing Centre (CSCS) under project ID a05 on Alps.
@@ A.1 Reaction Prediction (RxP)
Dataset Source USPTO-480K [1]
@@
The dataset for the RLSF training of this task was derived from the USPTO-480K [1] after removing the samples used in the SFT phase. 50K reactions were randomly chosen for the training set, and 500 reactions for the test set.
@@ A.2 IUPAC to SMILES (I2S)
Dataset Source PubChem (CID 2,000,001 to 3,000,000) [66]
@@ A.3 Reaction Naming
Dataset Source USPTO Reaction 1M; [84]
@@ A.4 Reaction Replacement
Dataset Source USPTO Reaction 1M; [84]
@@ A.5 Reaction Inversion
Dataset Source USPTO Reaction 1M; [84]
@@ A.6 Reaction True/False
Dataset Source USPTO Reaction 1M; [84]
@@ A.7 Chemical Formula Balancing Task (CeB)
Dataset Source The Perovskite Dataset [85]
@@ A.8 Conditional Material Generation (CMG)
Dataset Source Materials Project [86]
@@
Reward Function Design: The quality of the generated composition is measured by the metrics: validity, precision and novelty. Validity is assessed using SMACT [87] validity, which checks whether the generated composition adheres to fundamental chemical rules, such as charge neutrality. Precision measures the model’s ability to follow instructions and correctly include the specified elements. It is computed using the following equation:
@@
where $E_{pi}$ is the set of elements specified in the i-th prompt and $E_{qi}$ is the corresponding generated element [76]. The novelty of the generated composition was determined based on whether the composition was present within the materials project dataset or was previously generated by the model. Furthermore, to ensure the model provided its generated solution in a valid format, the reward function also checked that the generated composition was enclosed within the <material>…</material> tags and that the assigned space group number lies within the valid range of 1 to 230.
@@ A.9 Binary Compound Structure Relaxation Task (CrR)
Dataset Source Materials Project Trajectory (MPtrj) Dataset [88], Materials Project [86]
@@ B.1.1 Symbolic Competence Score benchmark
The Symbolic Competence Score benchmark measures the model’s latent capability to read and write correct chemical symbols. In this benchmark we focus particularly on SMILES, as organic chemistry spans a majority of our tasks. For this we collected 10000 valid SMILES from PubChem [66], such that no overlap exists with the MiST data. A second dataset is created with corrupted smiles based on these smiles, where corruptions are minimal, however render the smiles unvalid. The corruption procedure is specified in Algorithm 1. The algorithm removes a random subset of key structural grammar elements (ring/branch brackets and digits) from the SMILES string, producing broken or ambiguous strings. Corruption rate $\rho$ controls the proportion of removed elements, which for all our experiments has been set to 0.2.
@@ B.1.2 Chemical Competence Score benchmark
The Chemical Competence Score (CCS) evaluates a model’s latent ability to distinguish between chemically accurate and inaccurate factual statements. To construct this benchmark, we selected 1,000 samples from the test split of the SMolInstruct Molecule Description dataset [52], which was never used in all post-training stages. Each sample in the dataset consists of a brief description of an organic molecule. For example, one entry describes an acetamide as:
@@
For material data, we utilized Robocrystallographer [67] to generate  600 natural text descriptions for crystal structures from the Material Project [86]. Here is an example entry:
@@ D.1.1 FineWeb-Edu
The FineWeb-Edu can be found on Hugging Face (https://huggingface.co/datasets/HuggingFaceFW/fineweb-edu[89]. The subsets "CC-MAIN-2013-20" to "CC-MAIN-2024-10" were downloaded for a total of $\sim$6 TB, which represents roughly 1.3T tokens and 1.26B individual texts. Based on the representative subset "sample-10BT" (also downloaded), the text sources were computed by taking the base URL (from the dataset column "url"), then these sources were sorted from the most prevalent to the least. We manually labeled the most prevalent sources as "chemistry", "non-chemistry", or "undetermined". The goal was to label a source as "chemistry" only if nearly all the texts from that source are about chemistry. On the other hand, a source is classified as "non-chemistry" only if there is no mention of chemistry in all the texts from that source. When a source contains a mix, like a school website with chemistry texts and texts for other fields, the label used is "undetermined", and the source is not used. After this manual labeling, the texts from "sample-10BT" were classified based on the labeled sources. It led to a ground truth of approximately 10,000 "chemistry" texts and 50,000 "non-chemistry" texts (out of the $\sim$10M texts found in "sample-10BT"). Based on this ground truth, a custom non-ML classifier was built using the word frequencies in "chemistry" and "non-chemistry" texts. The texts were lemmatized before building word frequency vectors for the two classes using a simple processing script that replaces any non-standard character with a space, before splitting the strings by the spaces. A custom vocabulary was also built to store these lemmatized texts in a tokenized manner. Other lemmatization methods (such as Spacy or NLTK) were also tried, but did not lead to better results and were extremely expensive to use on the full FineWeb dataset (>>6 TB). After building the vocabulary and the word frequency vectors for the two classes, the formula below was applied to each FineWeb text to create an associated "chemistry score" (ranging from 0 to "infinity"). The frequencies of the lemma $k$ in chemistry texts and non-chemistry texts are denoted $f^{c}_{k}$ and $f^{n}_{k}$, respectively. The text chemistry score (TCS) is computed using the following equation:
@@ D.1.2 Academic paper extraction
An overview of our preprocessing pipeline is depicted as follows. Initially, we leveraged Nougat [90] and GROBID [91] libraries for converting PDF documents into textual formats. Nougat demonstrated superior performance in accurately transforming complex structures such as tables, formulae, bibliographic references, and figure captions into LaTeX-formatted text. Conversely, GROBID excelled at extracting plain textual content from PDFs. The output of the authors were merged with explicit tags assigned to each structural element: tables were encapsulated with [START_TABLE] and [END_TABLE], formulas marked by [START_FORMULA] and [END_FORMULA], bibliographic references enclosed within [START_BIBREF] and [END_BIBREF], and figure descriptions bracketed by [START_FIGURE] and [END_FIGURE]. Subsequently, this structured text was processed through the Chemical Data Extractor 2 [92], identifying candidate molecule entities along with their positional context within the text. To ensure high precision in entity identification, candidates were further validated using a custom-trained sentence transformer model, designed specifically to discern genuine molecular entities from contextual information. Validated molecular entities were then translated from their IUPAC nomenclature to SMILES notation using py2opsin, a Python interface for OPSIN [93]. In cases where OPSIN failed to yield a definitive conversion, entities were cross-referenced against PubChem [66]. Ultimately, during the pretraining phase alone, our model encountered approximately 800,000 unique chemical compounds along with their corresponding SMILES representations.
@@ D.1.3 PubChem
The first three million compounds from the PubChem database [66] (CID from 1 to 3,000,000) were dumped using the PUG REST API with batched requests in October 2024. Each record contains these columns (among others): CanonicalSMILES, IsomericSMILES, IUPACName, and InChI. Since the molecule canonicalization algorithm used in the PubChem database is not the same as the one used by RDKit, all the compounds were re-canonicalized. The canonical SMILES consistency was also ensured for each compound by computing four canonical SMILES for each molecule:
@@ D.2 Chemical Tasks Data sources
  • The USPTO-480K dataset [1] consists of approximately 480K organic reactions, divided into training and test splits.
@@ D.3 Material Tasks Data sources
  • A total of 1500 chemical formulas were selected from the Perovskite Dataset [85] to form the data set, and the data set was then enhanced by randomly masking individual stoichiometric coefficients within products or entire product compounds using [MASK].
@@
  • We selected 1000 samples from Materials Project [86] and extracted the constituent elements from each sample to create our dataset. For example, the compound TeO2 was decomposed into its constituent elements Te and O to form our training set.,
@@
  • We selected 2,000 binary compound crystal structures from the Materials Project [86] across the following categories: Intermetallics, Semiconductors, Oxides, Sulfides, Nitrides, Carbides, Hydrides, Halides, Borides, Silicides, Phosphides, Arsenides, Tellurides, and Selenides. And we applied perturbations to alter the positions of certain atoms and modify the cell parameters of these structures to form our training dataset.
@@ F.1 MiST
We conducted other experiments to evaluate our MiST model’s performance on other tasks and in comparison with strong baselines from the literature. In particular, we compare against NatureLM [76] and other general-purpose LLMS, on the task of SMILES to IUPAC and IUPAC to SMILES conversion. The results shown below put our MiST model (3B) on par with NatureLM 8B, while approaching the 8x7B MoE variant on IUPAC-to-SMILES conversion.

### v2

@@ 7 Limitations
While MiST demonstrates that targeted mid-stage pre-training can unlock chemical reasoning in a 3B-parameter model, several caveats remain. First, the SCS criterion works especially well in domains and tasks where the outcome validity is verifiable and easily corruptible, as is the case with chemistry and SMILES notation. We find such complications in the case of CMG, where CIF files are less-trivially corruptible, leading to worsened predictive power. Using a 100% valid notation, as is the case of SELFIES [selfies] for molecule generation tasks, can also compromise the value of SCS-like measures. Similar metrics can however be devised for biological sequential data, mathematical notation, etc. Second, the RLSF rewards we use focus on syntactic agreement with ground truth (e.g. exact SMILES or high Tanimoto similarity) and thus do not discourage chemically implausible or unsafe outputs, leaving open the possibility of reward hacking. Third, our evaluation suite—reaction prediction, IUPAC to SMILES translation, and conditional material generation, cover a narrow slice of chemistry; tasks that hinge on stereochemistry, kinetics, spectroscopy, or three-dimensional conformations remain unexplored. Finally, our pre-training corpus is dominated by small-molecule, chemical literature and patents, potentially biasing the model against macromolecular or bio-chemical domains. Addressing these limitations will be critical before SCS can be routinely used as a diagnostic metrics in other domains, and for MiST-style models to be relied upon as general scientific reasoning models.
@@ Acknowledgments
his work was supported as part of the Swiss AI Initiative by a grant from the Swiss National Supercomputing Centre (CSCS) under project ID a05 and a131 on Alps. AMB, SP, and PS acknowledge support from the NCCR Catalysis (grant number 225147), a National Centre of Competence in Research funded by the Swiss National Science Foundation. TX, BH, RZ also acknowledge resources provided by the UNSW Katana High Performance Computing cluster, and the National Computational Infrastructure (NCI Australia) under Project ID bz69. This work was supported by the Australian Renewable Energy Agency through the Australian Centre for Advanced Photovoltaics (ACAP). Responsibility for the views, information, or advice expressed herein is not accepted by the Australian Government.
@@ A.1 Reaction Prediction (RxP)
Dataset Source USPTO-480K [coley2019graph]
@@
The dataset for the RLSF training of this task was derived from the USPTO-480K [coley2019graph] after removing the samples used in the SFT phase. 50K reactions were randomly chosen for the training set, and 500 reactions for the test set.
@@ A.2 IUPAC to SMILES (I2S)
Dataset Source PubChem (CID 2,000,001 to 3,000,000) [pubchem_2025]
@@ A.3 Reaction Naming
Dataset Source USPTO Reaction 1M; [coley_prediction_2017]
@@ A.4 Reaction Replacement
Dataset Source USPTO Reaction 1M; [coley_prediction_2017]
@@ A.5 Reaction Inversion
Dataset Source USPTO Reaction 1M; [coley_prediction_2017]
@@ A.6 Reaction True/False
Dataset Source USPTO Reaction 1M; [coley_prediction_2017]
@@ A.7 Chemical Formula Balancing Task (CeB)
Dataset Source The Perovskite Dataset [Jacobsson2022FAIR]
@@ A.8 Conditional Material Generation (CMG)
Dataset Source Materials Project [matproj]
@@
Reward Function Design: The quality of the generated composition is measured by the metrics: validity, precision and novelty. Validity is assessed using SMACT [smact] validity, which checks whether the generated composition adheres to fundamental chemical rules, such as charge neutrality. Precision measures the model’s ability to follow instructions and correctly include the specified elements. It is computed using the following equation:
@@
where $E_{pi}$ is the set of elements specified in the i-th prompt and $E_{qi}$ is the corresponding generated element [naturelm]. The novelty of the generated composition was determined based on whether the composition was present within the materials project dataset or was previously generated by the model. Furthermore, to ensure the model provided its generated solution in a valid format, the reward function also checked that the generated composition was enclosed within the <material>…</material> tags and that the assigned space group number lies within the valid range of 1 to 230.
@@ A.9 Binary Compound Structure Relaxation Task (CrR)
Dataset Source Materials Project Trajectory (MPtrj) Dataset [Deng2023], Materials Project [matproj]
@@ B.1.1 Symbolic Competence Score benchmark
The Symbolic Competence Score benchmark measures the model’s latent capability to read and write correct chemical symbols. In this benchmark we focus particularly on SMILES, as organic chemistry spans a majority of our tasks. For this we collected 10000 valid SMILES from PubChem [pubchem_2025], such that no overlap exists with the MiST data. A second dataset is created with corrupted smiles based on these smiles, where corruptions are minimal, however render the smiles unvalid. The corruption procedure is specified in Algorithm 1. The algorithm removes a random subset of key structural grammar elements (ring/branch brackets and digits) from the SMILES string, producing broken or ambiguous strings. Corruption rate $\rho$ controls the proportion of removed elements, which for all our experiments has been set to 0.2.
@@ B.1.2 Chemical Competence Score benchmark
The Chemical Competence Score (CCS) evaluates a model’s latent ability to distinguish between chemically accurate and inaccurate factual statements. To construct this benchmark, we selected 1,000 samples from the test split of the SMolInstruct Molecule Description dataset [yu2024llasmol], which was never used in all post-training stages. Each sample in the dataset consists of a brief description of an organic molecule. For example, one entry describes an acetamide as:
@@
For material data, we utilized Robocrystallographer [ganose2019robocrystallographer] to generate  600 natural text descriptions for crystal structures from the Material Project [matproj]. Here is an example entry:
@@ D.1.1 FineWeb-Edu
The FineWeb-Edu can be found on Hugging Face (https://huggingface.co/datasets/HuggingFaceFW/fineweb-edu[penedo_fineweb_2024]. The subsets "CC-MAIN-2013-20" to "CC-MAIN-2024-10" were downloaded for a total of $\sim$6 TB, which represents roughly 1.3T tokens and 1.26B individual texts. Based on the representative subset "sample-10BT" (also downloaded), the text sources were computed by taking the base URL (from the dataset column "url"), then these sources were sorted from the most prevalent to the least. We manually labeled the most prevalent sources as "chemistry", "non-chemistry", or "undetermined". The goal was to label a source as "chemistry" only if nearly all the texts from that source are about chemistry. On the other hand, a source is classified as "non-chemistry" only if there is no mention of chemistry in all the texts from that source. When a source contains a mix, like a school website with chemistry texts and texts for other fields, the label used is "undetermined", and the source is not used. After this manual labeling, the texts from "sample-10BT" were classified based on the labeled sources. It led to a ground truth of approximately 10,000 "chemistry" texts and 50,000 "non-chemistry" texts (out of the $\sim$10M texts found in "sample-10BT"). Based on this ground truth, a custom non-ML classifier was built using the word frequencies in "chemistry" and "non-chemistry" texts. The texts were lemmatized before building word frequency vectors for the two classes using a simple processing script that replaces any non-standard character with a space, before splitting the strings by the spaces. A custom vocabulary was also built to store these lemmatized texts in a tokenized manner. Other lemmatization methods (such as Spacy or NLTK) were also tried, but did not lead to better results and were extremely expensive to use on the full FineWeb dataset (>>6 TB). After building the vocabulary and the word frequency vectors for the two classes, the formula below was applied to each FineWeb text to create an associated "chemistry score" (ranging from 0 to "infinity"). The frequencies of the lemma $k$ in chemistry texts and non-chemistry texts are denoted $f^{c}_{k}$ and $f^{n}_{k}$, respectively. The text chemistry score (TCS) is computed using the following equation:
@@ D.1.2 Academic paper extraction
An overview of our preprocessing pipeline is depicted as follows. Initially, we leveraged Nougat [nougat] and GROBID [grobid] libraries for converting PDF documents into textual formats. Nougat demonstrated superior performance in accurately transforming complex structures such as tables, formulae, bibliographic references, and figure captions into LaTeX-formatted text. Conversely, GROBID excelled at extracting plain textual content from PDFs. The output of the authors were merged with explicit tags assigned to each structural element: tables were encapsulated with [START_TABLE] and [END_TABLE], formulas marked by [START_FORMULA] and [END_FORMULA], bibliographic references enclosed within [START_BIBREF] and [END_BIBREF], and figure descriptions bracketed by [START_FIGURE] and [END_FIGURE]. Subsequently, this structured text was processed through the Chemical Data Extractor 2 [cde], identifying candidate molecule entities along with their positional context within the text. To ensure high precision in entity identification, candidates were further validated using a custom-trained sentence transformer model, designed specifically to discern genuine molecular entities from contextual information. Validated molecular entities were then translated from their IUPAC nomenclature to SMILES notation using py2opsin, a Python interface for OPSIN [opsin]. In cases where OPSIN failed to yield a definitive conversion, entities were cross-referenced against PubChem [pubchem_2025]. Ultimately, during the pretraining phase alone, our model encountered approximately 800,000 unique chemical compounds along with their corresponding SMILES representations.
@@ D.1.3 PubChem
The first three million compounds from the PubChem database [pubchem_2025] (CID from 1 to 3,000,000) were dumped using the PUG REST API with batched requests in October 2024. Each record contains these columns (among others): CanonicalSMILES, IsomericSMILES, IUPACName, and InChI. Since the molecule canonicalization algorithm used in the PubChem database is not the same as the one used by RDKit, all the compounds were re-canonicalized. The canonical SMILES consistency was also ensured for each compound by computing four canonical SMILES for each molecule:
@@ D.2 Chemical Tasks Data sources
  • The USPTO-480K dataset [coley2019graph] consists of approximately 480K organic reactions, divided into training and test splits.
@@ D.3 Material Tasks Data sources
  • A total of 1500 chemical formulas were selected from the Perovskite Dataset [Jacobsson2022FAIR] to form the data set, and the data set was then enhanced by randomly masking individual stoichiometric coefficients within products or entire product compounds using [MASK].
@@
  • We selected 1000 samples from Materials Project [matproj] and extracted the constituent elements from each sample to create our dataset. For example, the compound TeO2 was decomposed into its constituent elements Te and O to form our training set.,
@@
  • We selected 2,000 binary compound crystal structures from the Materials Project [matproj] across the following categories: Intermetallics, Semiconductors, Oxides, Sulfides, Nitrides, Carbides, Hydrides, Halides, Borides, Silicides, Phosphides, Arsenides, Tellurides, and Selenides. And we applied perturbations to alter the positions of certain atoms and modify the cell parameters of these structures to form our training dataset.
@@ F.1 MiST
We conducted other experiments to evaluate our MiST model’s performance on other tasks and in comparison with strong baselines from the literature. In particular, we compare against NatureLM [naturelm] and other general-purpose LLMS, on the task of SMILES to IUPAC and IUPAC to SMILES conversion. The results shown below put our MiST model (3B) on par with NatureLM 8B, while approaching the 8x7B MoE variant on IUPAC-to-SMILES conversion.